# Can Emulating Semantic Translation Help LLMs with Code Translation? A Study Based on Pseudocode

### v2

@@ 4.1. RQ1: Overall Effectiveness of Pseudocode-based Transitive Code Translation
Usefulness of Original Programs. The comparison between the pure pseudocode-based strategies with and without the original program (i.e., $\mathcal{P}$ vs. $\mathcal{PC}$) further reveals the need to include the original program in pseudocode-based code translation. Specifically, the pure pseudocode-based strategy with the original program as context ($\mathcal{PC}$) achieved an average pass@10 rate of 0.9270, 0.8357, and 0.6373 on the three levels’ solution programs, respectively, while the strategy without the original program ($\mathcal{P}$) only achieved pass@10 rates of 0.8634, 0.7392, and 0.5191 accordingly. The results demonstrate that the pseudocode generated by the translator LLM missed or misinterpreted details necessary for the correct generation of programs in the target PL in some cases. We investigate the concrete symptoms of this issue via case studies and discuss them in Section 4.5.
@@
Meanwhile, the helpfulness of including original programs becomes less obvious for the hybrid strategies (i.e., $\mathcal{D}\&\mathcal{P}$ vs. $\mathcal{D}\&\mathcal{PC}$). Specifically, the pass@10 rates of $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$ are generally comparable (with less than 0.01 difference in pass@10) in most comparisons. This indicates that the original program is less necessary to supplement pseudocode when using hybrid strategies, where direct translation may already provide access to referencing details in the original program for cases that highly rely on such details. The exceptional cases include GPT-4o-mini on hard-level tasks and Qwen7B on medium-level and hard-level tasks, where the hybrid strategy with the original program as context ($\mathcal{D}\&\mathcal{PC}$) achieved higher pass@10 rates. Therefore, it may still be beneficial to include the original program as a backup reference implementation of pseudocode when translating relatively complicated programs.
@@ 4.2. RQ2: Effectiveness Across Different Source and Target Programming Languages
Differences in syntax and semantics among PLs may cause varying translation difficulty across PL pairs (Pan et al., 2024; Macedo et al., 2025). In this RQ, we thus further study the effectiveness of the pseudocode-based translation strategies on different PL pairs. The comparison aims to (1) verify whether the helpfulness of pseudocode-based translation identified in RQ1 persists across different source and target PLs, and (2) explore if the pseudocode-based translation leads to significant improvements for certain PL pairs. The empirical findings are intended to guide developers in adopting pseudocode-based approaches for code translation involving different PLs. In this RQ, we focus on the improvement of the hybrid translation strategy with code ($\mathcal{D}\&\mathcal{PC}$), which exhibited generally optimal performance in RQ1, relative to the conventional direct translation strategy ($\mathcal{D}$).
(a) Qwen-32B (Easy Tasks)
(b) Qwen-32B (Med Tasks)
(c) Qwen-32B (Hard Tasks)
@@
(j) Qwen-7B (Easy Tasks)
(k) Qwen-7B (Med Tasks)
(l) Qwen-7B (Hard Tasks)
@@
Figure 4. Relative Improvement in Pass@10 Using “Direct + viaPseudocode (w/ code)” Strategy Compared to “Direct” Strategy for Code Translation. (The greener, the more improvement.)
@@
Overall Effectiveness. Figure 4 presents the relative improvement in pass@10 rates of the hybrid strategy $\mathcal{D}\&\mathcal{PC}$ over the baseline $\mathcal{D}$ across different PL pairs. Each row/column in a heatmap corresponds to a specific source/target PL. In general, we observed that the translation accuracy between almost all PL pairs improves (as indicated by positive values in figures) when using the $\mathcal{D}\&\mathcal{PC}$ strategy compared to $\mathcal{D}$. Across all 450 PL-pair-model-difficulty combinations, 147 (32.7%) combinations achieve even more than 10% improvement. The improvement is consistently observed across different LLMs and task difficulties, with more significant improvement on harder tasks and weaker LLMs, as observed in RQ1. The results confirm the generalizable benefits of adopting pseudocode-based translation among different PL pairs.
@@
Comparison Among PL Pairs. As shown by the different colors in heatmaps within Figure 4, the improvement varies across different PL pairs, which may be due to the differences across PL pairs and the varying training corpus of different PLs. Nevertheless, we observe consistent trends across different LLMs and task difficulties. Specifically, we first noticed that the improvement of code translation from flexible and lightweight PLs (e.g., Python and JavaScript) to stricter and more complex PLs (e.g., Rust and Go) is usually more significant than the improvement in the opposite direction, as shown by the deeper green colors near the right-bottom corner of each heatmap. For example, when conducting Python-to-Rust translation and JavaScript-to-Go translation on medium-level tasks, the improvement on Qwen32B is 13.64% and 8.96%, respectively; in comparison, it is only 5.54% and 4.05% in the opposite direction, respectively. We conjecture this is because transitive translation based on pseudocode helps guide LLMs to generate code fitting the requirements enforced in stricter and more complex PLs (e.g., Python and JavaScript are flexible with dynamic typing and rich syntax sugar, while C++ and Java require static type declaration and more rigid syntax, and Go and Rust further enforce more rules on variable usage and memory management).
@@
In addition, we observed that code translation to Rust (rightmost column in heatmaps) generally gains significant improvement, regardless of the source PL. We conjecture this may also result from the relatively lower resource of the Rust training corpus (Baltaji et al., 2025), particularly parallel code pairs between Rust and other PLs, which are vital to code-to-code translation (Ahmad et al., 2023) while limited in the practical training corpus. In comparison, there may be more natural language-Rust code pairs, such as Rust documentation and programming tutorials, which help LLMs learn to generate Rust code matching natural language descriptions. The pseudocode-based translation enables LLMs to generate Rust code in this manner, which may better align with the training corpus. Also, translating from Rust (top row in heatmaps) gains noticeable improvement in many trials, suggesting that pseudocode may help the understanding of original programs written in Rust. The results demonstrate the helpfulness of pseudocode in assisting code translation involving a low-resource PL.
@@
Message 3: Pseudocode-based translation generally improves code translation accuracy across most PL pairs. Besides, the improvement is more significant when translating from flexible and lightweight PLs to strict and complex PLs, and when translation involves the low-resource Rust. Practitioners are highly advised to consider pseudocode-based translation in these scenarios.
@@ 4.3. RQ3: Comparison with Programming Language as Transitive Intermediary
(a) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Python}}$ (Qwen-32B)
(b) vs. $\mathcal{D}\&\mathcal{PL}_{\text{JavaScript}}$ (Qwen-32B)
(c) vs. $\mathcal{D}\&\mathcal{PL}_{\text{C++}}$ (Qwen-32B)
(d) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Java}}$ (Qwen-32B)
(e) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Go}}$ (Qwen-32B)
(f) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Rust}}$ (Qwen-32B)
@@
Figure 5. Relative Improvement in Pass@10 Rates of Qwen-32B with Hybrid Strategy based on Pseudocode Compared to an Intermediate PL. (The greener, the more improvement. Red value means accuracy drop.)
@@
Figure 5 presents the comparison results of pseudocode-based translation ($\mathcal{D}\&\mathcal{PC}$) versus intermediate PL-based translation ($\mathcal{D}\&\mathcal{PL}_{L_{i}}$) on Qwen32B. The results show that pseudocode-based translations achieve higher code translation accuracy than intermediate PL-based translations in most cases. Specifically, among all 120 comparisons ($5$ source PLs $\times$ 44 target PLs $\times$ 66 transitive PLs), pseudocode-based translation wins in 109 comparisons (green cells), whereas it loses in only 11 comparisons (red cells) with marginal differences (less than 1%). Although certain intermediate PLs are more helpful than pseudocode for some source-target PL pairs (e.g., using JavaScript as the intermediate PL for Rust-to-Python, Go-to-C++, Go-to-Java, and C++-to-Java translation, Figure 5(b)), the pseudocode-based translations yield a generally more pronounced outperformance. The observed trends and conclusions are consistent across the other experimental LLMs (i.e., Qwen7B, Phi-4, GPT-4o-mini, and DSCoder). To avoid redundancy, we omit their detailed results here. These results are available in our artifact (Artifact of this paper, [n. d.]).
@@
Through case studies, we found that the advantage of pseudocode over PLs is its role as a general abstraction to bridge differences across diverse source-target PL pairs. Specifically, aligning with findings in (Macedo et al., 2025), we observed that a transitive PL is effective when it bridges the different features and APIs between the source and target PLs; meanwhile, different PL pairs may require varying transitive PLs. An inappropriate choice of intermediate PL may even increase the gap between the source and target PLs and introduce extra burdens in code translation, hindering the translation accuracy. Besides, an intermediate PL may also bring LLMs new burdens to satisfy its syntax and semantics, complicating the translation process. In contrast, flexible and descriptive pseudocode describes code intent and logic, offering a PL-agnostic abstraction that guides LLMs to focus on semantics instead of detailed PL-specific implementations, thereby offering better effectiveness in general. We also examined cases where pseudocode underperformed PLs in transitive translation and found that inferiority mainly stemmed from the missing essential details in the pseudocode (which will be discussed in Section 4.5, Limit-2). In comparison, programs in intermediate PLs often preserve the complete code semantics, facilitating the generation of a semantically consistent translation result. These findings highlight the advantages of pseudocode as a general effective intermediary, as well as the need for enhancing its quality to fully harness its effectiveness.
@@
Message 4: Pseudocode serves as a more general and effective intermediary than a specific PL in transitive code translation for most studied PL pairs. This is because pseudocode provides a PL-agnostic abstraction of code semantics that generally bridges differences across diverse source-target PL pairs, as an effective transitive PL acts for their fitting PL pairs.
@@ 4.4. RQ4: Effectiveness of Code Translation based on High-Quality Pseudocode
The pseudocode generated by the studied LLMs may introduce noise (e.g., incorrect or ambiguous descriptions of code intent and logic, which will be discussed in Section 4.5). Such low-quality pseudocode can mislead code translation. This RQ therefore examines whether better performance can be achieved when high-quality pseudocode is provided and identifies the bottlenecks in pseudocode-based code translation. Specifically, Wu et al. (2025) demonstrate that DeepSeek-R1 can help annotate human-written-like pseudocode with high accuracy and naturalness for concrete programs. Thus, in this RQ, we take DeepSeek-R1 as a high-quality pseudocode generator [^1] To avoid potential data contamination of DeepSeek-R1, whose knowledge cutoff date is unknown, we only adopt it to prepare high-quality pseudocode following Wu et al. (2025). We do not include it as a translator LLM to study.. Then, we evaluate the translation accuracy of the studied LLMs based on DeepSeek-R1-generated high-quality pseudocode. The investigation results help indicate the potential effectiveness of pseudocode-based translation when higher-quality pseudocode is available (e.g., written by humans or generated by enhanced pseudocode generation methods).
@@
Table 3. Pass@10 of Code Translation Based on Pseudocode Generated by Translator LLM (Self-Gen) and DeepSeek-R1 (R1-Generated).
@@
Self-Gen R1-Generated Self-Gen R1-Generated Self-Gen R1-Generated
@@
$\mathcal{D}\&\mathcal{PC}$ 0.9592 0.9750 (+1.64%) 0.8905 0.9149 (+2.74%) 0.7051 0.7481 (+6.10%)
(+xx%): performance improvement of a strategy based on R1-generated pseucode compared to LLM-translator-generated pseudocode.
Improvement with High-quality Pseudocode. Table 3 compares the pass@10 rates of different translation strategies based on high-quality pseudocode generated by DeepSeek-R1 and pseudocode generated by the studied LLMs themselves (i.e., the default setup studied in previous RQs), respectively. The results show that the code translation accuracy consistently improves across LLMs and strategies when using R1-generated pseudocode. Among all strategies, the pure pseudocode-based strategy ($\mathcal{P}$) benefits the most, with average improvements of 8.16%, 13.45%, and 22.15% on three difficulty levels, respectively, mitigating the underperformance relative to direct translation ($\mathcal{D}$) observed in RQ1. These results indicate pseudocode quality as a critical bottleneck for pseudocode-based translation. In addition, we observed that hybrid strategies ($\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$) continuously outperform single-approach strategies ($\mathcal{D}$, $\mathcal{P}$, $\mathcal{PC}$), indicating that combining direct and pseudocode-based translation remains beneficial even with higher-quality pseudocode. With R1-generated high-quality pseudocode, the best pass@10 scores of all LLMs with optimal strategy improved to 0.9646-0.9835, 0.8861–0.9512, and 0.6747–0.8286 across the three difficulty levels, demonstrating the promising potential of pseudocode-based code translation.
Message 5: The quality of pseudocode hinders accurate pseudocode-based translation of the studied LLMs. Higher-quality pseudocode can consistently improve the code translation accuracy of studied LLMs across translation strategies and task difficulties.
LLM-wise Comparison. The performance comparison among LLMs further demonstrates the remaining limitations of LLMs in both code understanding and generation during pseudocode-based code translation. Specifically, we observed that the performance improvement brought by R1-generated pseudocode is more significant for weaker LLMs. For example, the improvement of all strategies on Qwen7B ranges from 2.31% to 64.03%, while it is only 0.61% to 8.98% on Qwen32B, indicating that higher-quality pseudocode is more beneficial for weaker LLMs. This demonstrates the weakness of weaker LLMs in understanding and summarizing the code intent and logic of original programs. In addition, with R1-generated pseudocode, the performance gap between weaker LLMs and more powerful LLMs remains. For example, with R1-generated pseudocode, the pass@10 rates of Qwen7B and DSCoder with $\mathcal{D}\&\mathcal{PC}$ are only 0.6747 and 0.7073 on hard tasks, respectively, which are much lower than 0.8286 of Qwen32B. The results indicate that even given the same code intent and logic (i.e., pseudocode), the weaker LLMs are still less capable in code translation than the more powerful LLMs, indicating that the code implementation capability also hinders the effectiveness of pseudocode-based code translation on weaker LLMs.
Message 6: Both the code understanding capability (indicated by the quality of the generated pseudocode) and the code implementation capability of the studied LLMs hinder their performance in code translations.

### v3

@@ 4.1. RQ1: Overall Effectiveness of Pseudocode-based Code Translation
Usefulness of Original Programs. The comparison between the pure pseudocode-based strategies with and without the original program (i.e., $\mathcal{P}$ vs. $\mathcal{PC}$) further reveals the need to include the original program in pseudocode-based code translation. Specifically, the pure pseudocode-based strategy with the original program as context ($\mathcal{PC}$) achieved an average pass@10 rate of 0.9270, 0.8357, and 0.6373 on the three levels’ solution programs, respectively, while the strategy without the original program ($\mathcal{P}$) only achieved pass@10 rates of 0.8634, 0.7392, and 0.5191 accordingly. The results suggest that the pseudocode generated by the translator LLM missed or misinterpreted details necessary for the correct generation of programs in the target PL in some cases. We investigate the concrete symptoms of this issue via case studies and discuss them in Section 4.5.
@@
Meanwhile, the helpfulness of including original programs becomes less obvious for the hybrid strategies (i.e., $\mathcal{D}\&\mathcal{P}$ vs. $\mathcal{D}\&\mathcal{PC}$). Specifically, the pass@10 rates of $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$ are generally comparable (with less than 0.01 difference in pass@10) in most comparisons. This suggests that the original program is less necessary to supplement pseudocode when using hybrid strategies, as direct translation may already provide access to referencing details in the original program for translation tasks that heavily rely on such details, further confirming the complementary strengths of both approaches. The exceptional cases include GPT-4o-mini on hard-level tasks and Qwen7B on medium-level and hard-level tasks, where the hybrid strategy with the original program as context ($\mathcal{D}\&\mathcal{PC}$) achieved obviously higher pass@10 rates. Therefore, it is generally beneficial to include the original program as a backup reference implementation of pseudocode when translating relatively complicated programs.
@@ 4.2. RQ2: Effectiveness Across Different Source and Target Programming Languages
Differences in syntax and semantics among PLs may cause varying translation difficulty across PL pairs (Pan et al., 2024; Macedo et al., 2025). In this RQ, we thus further compare the effectiveness of the pseudocode-based translation strategies on different source-target PL pairs. The comparison aims to (1) verify whether the helpfulness of pseudocode-based translation identified in RQ1 persists across different source and target PLs, and (2) explore if the pseudocode-based translation leads to significant improvements for certain PL pairs. The empirical findings are intended to guide developers in adopting pseudocode-based approaches for code translation involving different PLs. In this RQ, we focus on the improvement of the hybrid translation strategy with context ($\mathcal{D}\&\mathcal{PC}$), which showed generally optimal performance in RQ1, relative to the conventional direct translation strategy ($\mathcal{D}$).
(a) Qwen32B (Easy Tasks)
(b) Qwen32B (Med Tasks)
(c) Qwen32B (Hard Tasks)
@@
(j) Qwen7B (Easy Tasks)
(k) Qwen7B (Med Tasks)
(l) Qwen7B (Hard Tasks)
@@
Figure 5. Relative Improvement in Pass@10 Using Hybrid Strategy $\mathcal{D}\&\mathcal{PC}$ Compared with Direct Translation Strategy $\mathcal{D}$ Across Different Source-Target Programming Languages. (The greener, the more improvement.)
@@
Overall Effectiveness. Figure 5 presents the relative improvement in pass@10 rates of the hybrid strategy $\mathcal{D}\&\mathcal{PC}$ over the baseline $\mathcal{D}$ across different PL pairs. Each row/column in a heatmap corresponds to a specific source/target PL. In general, we observed that the translation accuracy between almost all PL pairs improves (as indicated by positive values in figures) when using the $\mathcal{D}\&\mathcal{PC}$ strategy compared to $\mathcal{D}$. Across all 450 PL-pair-model-difficulty combinations, 147 (32.7%) combinations achieve even more than 10% improvement. The improvement is consistently observed across different LLMs and task difficulties, with more significant improvement on harder tasks and weaker LLMs, similar to trends observed in RQ1. The results confirm the generalizable benefits of adopting pseudocode-based translation among different PL pairs.
@@
Comparison Among PL Pairs. As shown by the different colors in heatmaps within Figure 5, the improvement varies across different PL pairs, which may be due to the differences across PL pairs and the varying training corpus of different PLs. Nevertheless, we observe several consistent trends across different LLMs and task difficulties. Specifically, we first noticed that the improvement of code translation from flexible and lightweight PLs (e.g., Python and JavaScript) to stricter and more complex PLs (e.g., Rust and Go) is often more significant than the improvement in the opposite direction, as shown by more dark green colors near the right-bottom corner of each heatmap. For example, when conducting Python-to-Rust translation and JavaScript-to-Go translation on medium-level tasks, the improvement on Qwen32B is 13.64% and 8.96%, respectively; in comparison, it is only 5.54% and 4.05% in the opposite direction, respectively. We conjecture this is because transitive translation based on pseudocode helps guide LLMs to generate code fitting the requirements enforced in stricter and more complex PLs (e.g., Python and JavaScript are flexible with dynamic typing and rich syntax sugar, while C++ and Java require static type declaration and more rigid syntax, and Go and Rust further enforce more rules on variable usage and memory management).
@@
In addition, we observed that code translation to Rust (rightmost column in heatmaps) generally gains significant improvement, regardless of the source PL. We conjecture this may also result from the relatively lower resource of the Rust training corpus (Baltaji et al., 2025), particularly parallel code pairs between Rust and other PLs, which are vital to code-to-code translation (Ahmad et al., 2023) while limited in the practical training corpus. In comparison, there may be more natural language-Rust code pairs, such as Rust documentation and programming tutorials, which help LLMs learn to generate Rust code matching natural language descriptions. The pseudocode-based translation enables LLMs to generate Rust code in this manner, which may better align with the training corpus. Also, translating from Rust (top row in heatmaps) gains noticeable improvement in many trials, suggesting that pseudocode also helps the understanding of original programs written in Rust. The results demonstrate the helpfulness of pseudocode in assisting code translation involving a low-training-resource PL.
@@
Message 3: Pseudocode-based translation generally improves code translation accuracy across most PL pairs. In addition, improvement is more significant when translating from flexible and lightweight PLs to strict and complex PLs, and when dealing with the low-training-resource Rust. Practitioners are highly advised to consider pseudocode-based translation in these scenarios.
@@ 4.3. RQ3: Comparison with Programming Language as Transitive Intermediary
(a) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Python}}$ (Qwen32B)
(b) vs. $\mathcal{D}\&\mathcal{PL}_{\text{JavaScript}}$ (Qwen32B)
(c) vs. $\mathcal{D}\&\mathcal{PL}_{\text{C++}}$ (Qwen32B)
(d) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Java}}$ (Qwen32B)
(e) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Go}}$ (Qwen32B)
(f) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Rust}}$ (Qwen32B)
@@
Figure 6. Relative Improvement in Pass@10 Rates of Qwen32B Using Hybrid Strategy Based on Pseudocode Compared to Using Hybrid Strategy Based on an Intermediate PL. (The greener, the more improvement. Red value means performance drop.)
@@
Figure 6 presents the comparison results of pseudocode-based translation ($\mathcal{D}\&\mathcal{PC}$) versus intermediate PL-based translation ($\mathcal{D}\&\mathcal{PL}_{L_{i}}$) on Qwen32B. The results show that pseudocode-based translations achieve higher code translation accuracy than intermediate PL-based translations in most situations. Specifically, among all 120 comparisons ($5$ source PLs $\times$ 44 target PLs $\times$ 66 transitive PLs), pseudocode-based translation wins in 109 comparisons (green cells), whereas it loses in only 11 comparisons (red cells) with marginal differences (less than 1%). Although certain intermediate PLs are more helpful than pseudocode for some source-target PL pairs (e.g., using JavaScript as the intermediate PL for Rust-to-Python, Go-to-C++, Go-to-Java, and C++-to-Java translation, Figure 6(b)), the pseudocode-based translations yield a generally more pronounced outperformance. The observed trends and conclusions are consistent across the other studied LLMs (i.e., Qwen7B, Phi-4, GPT-4o-mini, and DSCoder). To avoid redundancy, we omit their detailed results here. These results are available in our artifact (Artifact of this paper, ).
@@
Through case studies, we found that the advantage of pseudocode over PLs is its role as a general abstraction to bridge differences across diverse source-target PL pairs. Specifically, aligning with findings in (Macedo et al., 2025), we observed that a transitive PL is effective when it bridges the different features and APIs between the source and target PLs; meanwhile, different PL pairs may require varying transitive PLs. An inappropriate choice of intermediate PL may even increase the gap between the source and target PLs and introduce extra burdens in code translation, hindering the translation accuracy. In addition, an intermediate PL may also bring LLMs new burdens to satisfy its syntax and semantics, complicating the translation process. In contrast, flexible and descriptive pseudocode presents code intent and logic, offering a PL-agnostic abstraction that guides LLMs to focus on semantics instead of detailed PL-specific implementations, thereby offering better effectiveness in general. We also examined cases where pseudocode underperformed PLs in transitive translation and found that inferiority mainly stemmed from missing essential details in the pseudocode (which will be discussed in Section 4.5.2, Limit-2). In comparison, programs in effective intermediate PLs often preserve the complete code semantics, facilitating the generation of a semantically preserving translation result. These findings highlight the advantages of pseudocode as a general effective intermediary, as well as the need for enhancing its quality to fully harness its effectiveness.
@@
Message 4: Pseudocode serves as a more general and effective intermediary than a specific PL in transitive code translation for most studied PL pairs. This is because pseudocode provides a PL-agnostic abstraction of code semantics that generally bridges differences across diverse source-target PL pairs, as an effective transitive PL does for suitable PL pairs. In development, practitioners can prioritize pseudocode as a universal bridge for translation across various PLs.
@@ 4.4. RQ4: Effectiveness of Code Translation based on High-Quality Pseudocode
The pseudocode generated by the studied LLMs may introduce noise (e.g., incorrect or ambiguous descriptions of code intent and logic, which will be discussed in Section 4.5). Such low-quality pseudocode can mislead code translation. This RQ therefore examines whether better performance can be achieved when high-quality pseudocode is available, as well as identifies the bottlenecks in pseudocode-based code translation. Specifically, DeepSeek-R1 (DeepSeek-AI, 2025) is found capable of helping annotate human-written-like pseudocode with high accuracy and naturalness for concrete programs (Wu et al., 2025). Therefore, in this RQ, we take DeepSeek-R1 as a high-quality pseudocode generator [^2] To avoid potential data contamination of DeepSeek-R1, whose knowledge cutoff date is unknown, we only adopt it to prepare high-quality pseudocode following Wu et al. (2025). We do not include it as a translator LLM to study. and evaluate the translation accuracy of the studied LLMs based on DeepSeek-R1-generated high-quality pseudocode. The investigation results help indicate the potential effectiveness of pseudocode-based translation when higher-quality pseudocode is available (e.g., written by humans or generated by enhanced pseudocode generation methods).
@@
Table 4. Pass@10 of Round-Trip Code Regeneration Based on Pseudocode Generated by Translator LLM (Self-Gen) and DeepSeek-R1 (R1-Gen).
LLM Easy Tasks Medium Tasks Hard Tasks
LLM-Gen R1-Gen LLM-Gen R1-Gen LLM-Gen R1-Gen
Qwen32B 0.9516 0.9670 (+1.62%) 0.8976 0.9042 (+0.74%) 0.7352 0.8050 (+9.49%)
Phi-4 0.8889 0.9425 (+6.03%) 0.8010 0.8365 (+4.43%) 0.6738 0.6773 (+0.52%)
GPT-4o-mini 0.8865 0.9583 (+8.10%) 0.7583 0.8646 (+14.02%) 0.5532 0.6986 (+26.28%)
Qwen7B 0.8008 0.9267 (+15.72%) 0.6413 0.8156 (+27.18%) 0.3865 0.5567 (+44.04%)
DS-Coder 0.8774 0.9353 (+6.60%) 0.7833 0.8188 (+4.53%) 0.5875 0.6525 (+11.06%)
(+x.xx%): performance improvement of a strategy based on R1-generated pseudocode compared to translator-LLM-generated pseudocode.
Quality Check of R1-Generated Pseudocode: Before comparing the code translation accuracy based on different pseudocode, we verify the assumed higher quality of pseudocode generated by DeepSeek-R1 over the studied LLMs using round-trip validation, an effective method to validating quality of semantic representations like code summaries (Allamanis et al., 2024), natural language translations (Aiken and Park, 2010), and pseudocode (Wu et al., 2025). Specifically, we use each studied LLM to regenerate the original program based on the pseudocode generated by itself and DeepSeek-R1, respectively, following the prompt in Figure 3(b) with the original PL. The regeneration pass rates can indicate the pseudocode quality, since higher-quality pseudocode should better describe code intent and logic of original programs, thus facilitating regeneration. As shown in Table 4, the pass rates based on R1-generated pseudocode are consistently higher than translator-generated pseudocode, suggesting the higher quality of R1-generated pseudocode in describing the code intent and logic of original programs.
Improvement with High-quality Pseudocode. Table 5 compares the pass@10 rates of different translation strategies based on high-quality pseudocode generated by DeepSeek-R1, and pseudocode generated by the studied LLMs themselves (i.e., the default setup studied in previous RQs), respectively. The results demonstrate that the code translation accuracy consistently improves across LLMs and strategies after switching to DeepSeek-R1-generated pseudocode. Among all strategies, the pure pseudocode-based strategy ($\mathcal{P}$) benefited the most, with average improvements of 8.16%, 13.45%, and 22.15% on three difficulty levels, respectively, mitigating the underperformance relative to direct translation ($\mathcal{D}$) observed in RQ1. These results indicate pseudocode quality as a critical bottleneck for pseudocode-based translation. In addition, we observed that hybrid strategies ($\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$) continuously outperformed single-approach strategies ($\mathcal{D}$, $\mathcal{P}$, $\mathcal{PC}$), indicating that combining direct and pseudocode-based translation remains beneficial even with higher-quality pseudocode. With DeepSeek-R1-generated high-quality pseudocode, the best pass@10 scores of all LLMs with optimal strategy improved to 0.9646-0.9835, 0.8861–0.9512, and 0.6747–0.8286 across three difficulty levels, demonstrating the promising potential of pseudocode-based code translation.
Table 5. Pass@10 of Code Translation Based on Pseudocode Generated by Translator LLM (Self-Gen) and DeepSeek-R1 (R1-Gen).
@@
Self-Gen R1-Gen Self-Gen R1-Gen Self-Gen R1-Gen
@@
$\mathcal{D}\&\mathcal{PC}$ 0.9592 0.9750 (+1.64%) 0.8905 0.9149 (+2.74%) 0.7052 0.7481 (+6.09%)
(+x.xx%): performance improvement of a strategy based on R1-generated pseudocode compared to translator-LLM-generated pseudocode.
Message 5: The quality of pseudocode hinders accurate pseudocode-based code translation of the studied LLMs. Higher-quality pseudocode can consistently improve the code translation accuracy of studied LLMs across translation strategies and task difficulties. Practitioners may also consider human-in-the-loop to prepare high-quality pseudocode or revise the generated pseudocode, which will help guide LLMs to generate more accurate translation results through pseudocode-based code translation.
LLM-wise Comparison. The performance comparison among LLMs further demonstrates the limitations of LLMs in both code understanding and generation during pseudocode-based code translation. Specifically, we observed that the performance improvement brought by DeepSeek-R1-generated pseudocode is more significant for weaker LLMs. For example, the improvement of all strategies on Qwen7B ranges from 2.31% to 64.03%, while it is only 0.61% to 8.98% on Qwen32B, indicating that higher-quality pseudocode is more beneficial for weaker LLMs. This demonstrates the weakness of weaker LLMs in understanding and interpreting the code intent and logic of original programs. In addition, with DeepSeek-R1-generated pseudocode, the performance gap between weaker LLMs and more powerful LLMs remains. For example, with R1-generated pseudocode, the pass@10 rates of Qwen7B and DSCoder with $\mathcal{D}\&\mathcal{PC}$ are only 0.6747 and 0.7073 on hard tasks, respectively, which are much lower than 0.8286 of Qwen32B. The results indicate that even given the same code intent and logic (i.e., pseudocode), the weaker LLMs are still less capable in code translation than the more powerful LLMs, indicating that the code implementation capability also hinders the effectiveness of pseudocode-based code translation on weaker LLMs.
Message 6: Both code understanding capability (indicated by the quality of the interpreted pseudocode) and code implementation capability (based on pseudocode) of the studied LLMs hinder their performance in code translation tasks. Further advancements in generating high-quality pseudocode from programs and implementing accurate code from pseudocode are necessary to fully unleash the potential of pseudocode-based code translation.